A body rolled on the surface with a speed of 10m/s comes to rest after covering a distance of 50m. Find the coefficient of friction.g=10m/s.s
 

Grade:11

1 Answers

Arun
5 years ago

Average speed will be 10/2 = 5m/s. To cover 50m at 5m/s will take 10 sec. Deceleration will be 10/10 = 1m/s/s.

F = Ma

F = W/g (1) where W = weight of body. F = µW. where µ = coef of frict.

µW = W/g

µ = i/g = .1